# PrisonBreak: Jailbreaking Large Language Models with Fewer Than Twenty-Five Targeted Bit-flips

### v1

@@ 9.2 Behavioral Analysis
For an in-depth understanding of how our attack works, we analyze at the activation level focusing on two aspects: directional alignment between the activations produced by different attacks and a decomposition of them.
@@
Mechanistic analysis. Recent work [39] showed a refusal feature in language models, which determines whether the model will deny a harmful query. The refusal direction is the average difference in activations computed for the last token of harmful and benign queries (e.g., queries from HarmBench and Alpaca), and captures how the model differentiates between the two. We first compute the refusal direction for clean models following the procedure outlined by Arditi et al. [39]. We then measure the directional agreement between the last token activations of each attack with the refusal direction on data from the HarmBench test set, using cosine similarity. If the similarity is higher, it means the model is expressing the refusal feature, which theoretically should increase the probability of a refusal.
@@
In the leftmost plot in Figure 4, we compare the agreement across different attacks on Vicuna-13B-v1.5. Prompt-based attacks like GCG show less agreement, by reducing a model’s focus on harmful inputs and effectively bypassing refusal by limiting its expression [39]. Ortho removes the direction entirely (indicated by the constant cosine similarity of zero), suppressing a model’s ability to refuse. Interestingly, our attack reduces the expression of refusal comparatively less than the others. PrisonBreak shows much higher cosine similarities that closely align with the clean model. This indicates a strong presence of the refusal feature, which is surprising given that we show models jailbroken by PrisonBreak refuse at a lower rate. It does not seem to control the refusal direction to jailbreak models, thus functioning differently than Ortho.
Figure 4: Behavorial analysis results. For each attack, the (left) cosine similarity of HarmBench activations and the refusal direction on all layers of Vicuna-13b-v1.5, and the UMAP visualization of activations from Alpaca (middle) and HarmBench (right) from the last layer of Qwen2-7b.
@@
Decomposing activations. To further understand activation differences in jailbroken models, we decompose them using a dimensionality reduction technique, UMAP [64]. We draw 100 samples randomly-chosen from Alpaca and HarmBench, and for each attack, obtain the activation at the last token position. We run UMAP on activations collected from the 200 samples, and reduce them into 2 components [^3] Note that this approach compares activations from a clean model and those compromised by Ortho and ours. We empirically find that the compromised models are close to the clean model in the loss space, indicating that their activations will not differ drastically..
@@
The two remaining figures in Figure 4 show our visualization results on Qwen2-7B. The visualizations are from the last layer, where activations are most distinct and closest to the model’s output (see Appendix B for visualizations of more layers). The clean model effectively distinguishes between benign and harmful prompts, shown by the two distinct clusters. It is possible that the cluster of harmful activations represents a region in the activation space where the model has learned to refuse. Prompt-based attacks like GCG produces activations that are similar for both harmful and benign inputs. This confirms that the adversarial suffix induces a distribution shift in the inputs, bypassing refusal by producing activations that differ from what the model should identify. Similarly for Ortho, most activations from harmful inputs are very close to benign ones, and also align with the benign activations of the clean model. It suggests that Ortho suppresses the refusal features, causing the jailbroken model to respond to all inputs. Our attack shows a different decomposition. The cluster for benign inputs overlaps significantly with that from the clean model, but that of harmful inputs forms a distinct cluster. This indicates that our attack preserves a model’s generation capabilities but does not fully-suppress the refusal features like Ortho, as shown in our mechanistic analysis. PrisonBreak may instead amplify activations associated with responding.
@@ 10 Potential Countermeasures
Many defenses have been proposed to mitigate Rowhammer. This includes hardware-level defenses that employ row-refreshing and swapping strategies [65, 66, 67, 68, 69, 15, 70] or error correcting codes (ECC) [71, 72]. However, row-refreshing and swapping incur significant area and latency costs, and ECC have shown to be bypassable with advanced Rowhammer attacks [73, 74, 45]. System-level defenses, on the other hand, modify memory allocation protocols [75, 76, 77, 21], but they also impose costs as they require changes to memory allocators and specialized memory layouts, which introduce substantial memory and computational overhead. The limitations and overhead have restricted their adoption in real-world infrastructure.
Model-level defenses mitigate the impact of bitwise corruptions in parameters. They generally fall into two categories: detection and prevention. Prior work on detection focuses on identifying changes to weight values [78, 79, 80, 81] so that bitwise errors can be removed by reloading the target model. However, these defenses add high overhead by defending the entire model [81, 79] or only monitor parameters critical to untargeted classification [80, 78].
A separate line of work prevents changes to weight values [82, 83, 60, 84, 85]. One approach is to use memory representations that limit changes in weight values under bitwise corruption, such as quantization/binarization of model parameters [83, 60]. Recent work [82] instead proposes an architectural modification that introduces internal classifiers, allowing the model to stop propagating abnormal activations caused by bit-flips. While shown effective, these approaches require (re-)training of a model, which incurs substantial computational costs, especially for large-scale language models. Moreover, adapting architectural modifications designed for computer vision models to Transformers requires non-trivial effort [86, 87].
@@
We examine prevention mechanisms that do not greatly increase computational costs. We evaluate PrisonBreak against two popular defenses: Ranger [85], which clamps activations during inference, and quantized-weight reconstruction [84] which bounds weights within ranges determined by locally-computed averages. These defenses leverage the insight that bit-flip attacks tend to cause large changes to parameter values. For Ranger, we compute activation bounds using 20% of Alpaca data (following the original studies) and clip the activations from all Transformer-block outputs. We use a straightforward adaptation of the defense by Li et al. [84] that clips the weights of each layer to their original min. and max. values before each generation.
Defense Models None Ranger Weight Clip Weight Clip + Stealthy Vicuna-7b-v1.3 94.34 94.34 12.56 67.92 Llama2-7b 86.79 86.16 1.26 64.15 Qwen2-1.5b 80.50 81.13 15.09 78.62 Average 87.21 87.21 9.64 70.23 Time (ms) N/A 269.13 3.4 3.4
Table 8: Defense results. The ASR of PrisonBreak on model-level defenses. The average per-query computational overhead of each defense (Time) is reported in milliseconds using a single Nvidia A40 GPU.
Table 8 summarizes our results. Ranger is ineffective, showing no reduction in ASR because the massive activations used by language models [88] result in clamping bounds that are not exceeded despite large weight changes. Ranger also adds significant overhead, as activation clamping is required at each forward pass. In contrast, weight clipping effectively reduces ASR from 87.2% to 9.6%, close to the 9.4% level of clean models, with minimal computational overhead (3.4ms) since it only needs to be applied once per response. We also examine an adaptive adversary (Weight Clip + Stealthy) who selectively inflicts bit-flips that keep parameter changes within specified layer-wise bounds. The attacker regains effectiveness, achieving an ASR of 70.2%.
@@ 11 Conclusion
We introduce PrisonBreak, a new attack on language models where an adversary induces bit-flips at specific memory locations to jailbreak them in runtime. To this end, we present BitFinder, a computationally-efficient offline procedure for identifying target bits, and PrecisionHammer, which exploits Rowhammer attacks to flip those target bits. We demonstrate against 8 language models publicly available on the Internet that the attacker can PrisonBreak these models with as few as 5 and up to 25 bit-flips at most. Our jailbroken models generate high-quality responses to harmful queries with an attack success of 81–94% while preserving their utility. We also demonstrate that PrecisionHammer reliably jailbreaks all evaluated models on systems targeted by existing bit-flip attacks. Furthermore, we show that our hardware-aware search algorithm can achieve successful jailbreaking on these models, even on highly Rowhammer-secure systems. We characterize the mechanisms behind our attack, and show PrisonBreak operates differently than existing jailbreaking attacks. We also discuss the effectiveness of potential countermeasures to reduce the risk.
@@ Appendix A Experimental Setup in Detail
Here, we describe our experimental setup in detail. We implement our attack using Python v3.10.13 and PyTorch v2.4.1, which supports CUDA 12.1 for GPU usage. All language models and datasets used in our work are open-source and available on HuggingFace [57] or their respective repositories. Information regarding the DRAM modules used in our study is available from the original work [22]. We run BitFinder and our evaluation on a machine with an Intel Xeon Processor with 48 cores, 64GB memory, and 8 Nvidia A40 GPUs. For our fault injection simulations, we use an Intel Xeon Processor with 40 cores and 128GB memory.
Appendix B Activation Decomposition
@@
Figure 5 shows the decomposed activations from Alpaca and HarmBench for three additional layers (4, 12, and 20) using the procedure detailed in §9.2. In earlier layers, we find that benign and harmful activations are not clustered apart in the clean model, as it is likely focusing on lower-level information [63] (e.g., syntax). Ortho and our attack produce activations that are virtually identical to the clean model, because the parameter perturbations are minimal (in our case, we do not flip bits before or in layer 4). GCG, however, has already formed a distinct cluster, showing that the adversarial suffix has an immediate impact. In the middle layers, we see clusters forming between benign and harmful activations for all methods but GCG. Interestingly, the harmful activations from PrisonBreak and Ortho are clustered tightly together but separate from the clean model; the attack’s act on activations similarly here, despite forming distinct clusters in later layers. By layer 20, we observe a similar clustering to layer 12 except that the harmful activations from our attack and Ortho have separated. Many bit flips made by PrisonBreak are concentrated in these middle–final layers, evidently inducing a unique shift in the activation space.
Appendix C Results on tinyBenchmarks
Models tinyBenchmarks Tasks Arc GSM8k Hella- Swag MMLU Truth- FulQA Wino- Grande Vicuna-7B-v1.3 -0.58 +0.38 -1.16 -0.54 -1.98 -4.08 Vicuna-7B-v1.5 +0.03 -3.04 -0.97 -0.61 -7.92 -2.57 Vicuna-13B-v1.5 -1.34 -4.60 -4.98 -0.84 -5.90 -6.87 Llama2-7B +0.88 -4.54 -4.51 -1.77 -4.14 +2.58 Llama2-13B -5.68 -12.03 -4.53 -1.14 -0.53 -1.09 Llama3-8B -0.55 -2.40 -0.50 +0.50 +0.07 +3.22 Qwen2-1.5B -2.26 -11.44 -1.42 -1.25 -2.88 -5.32 Qwen2-7B +0.49 -3.67 -0.42 -2.47 +0.08 -0.04 Average -1.13 -5.12 -2.31 -1.02 -2.90 -1.77
@@ Appendix C Results on tinyBenchmarks
Table 9: Full tinyBenchmarks results. The accuracy change from PrisonBreak on each tinyBenchmark task for all models.
@@
Table 9 shows our full evaluation results for PrisonBreak on tinyBenchmarks. Notably, the change in performance differs across tasks. Models on average perform the worst on GSM8k—a benchmark consisting of grade school math problems—with an accuracy reduction of 5.12%. This may arise from the lack of these problems in the dataset we use to compute the utility score. The second most affected task is TruthfulQA with an average accuracy reduction of 2.90%. This fact is less surprising as jailbroken models are more likely to be deceptive [38, 39]. Other tasks exhibit smaller drops in accuracy; in general, we find that PrisonBreak preserves utility on benign tasks.
Figure 6: The ASR and Acc. across proxy dataset size, number of critical weights, utility score threshold, and number of bit-flips in Vicuna-7b-v1.3.
Appendix D Impact of Attack Configurations
In this section we vary the attack configuration choices for an additional model: Vicuna-7b-v1.3. Results are presented in Figure 6. For the proxy dataset size, using just 5 samples is sufficient to achieve a high ASR (87.42%). We also find that increasing the number of candidates from 100–150 surprisingly lowers ASR from 94.34–91.82%. These results differ from §7.4, where we found that a dataset size of $5$ is ineffective and setting $k=150$ substantially improved ASR on Llama2-7b. It is likely that because lesser-aligned models like Vicuna-7b-v1.3 are easier to jailbreak, PrisonBreak requires less data and fewer candidates. We observe largely similar results for the utility score threshold, but find the most aggressive value ($0.001$) less effective, as increasing it to $0.01$ raises ASR from 85.53–94.34% with negligible impact on Acc. We find that the tighter threshold required the analysis of substantially more candidates, meaning the selected bit-flips minimized the jailbreaking objective less effectively. Like §7.4, a manual analysis of the harmful generations finds that higher thresholds ($>0.01$) result in lower fidelity outputs regardless of the reported Acc. on tinyBenchmarks. Finally, the rightmost figure shows that only 5 bit-flips are needed to achieve the best ASR, after which it saturates; confirming our initial findings that lesser-aligned models require fewer bit-flips to jailbreak.

### v2

@@ 8.2 Behavioral Analysis
To gain a deeper understanding of how our attack works, we analyze directional alignment between the activations produced by different attacks and their decomposition.
@@
Mechanistic analysis. Recent work [34] identified a refusal feature in language models that determines if the model will deny a harmful query. This feature—the refusal direction—is computed as the average difference in activations computed for the last token of harmful (HarmBench) and benign queries (Alpaca). We compute the refusal direction for clean models following the procedure in Arditi et al. [34]. We then measure the directional agreement between the last token activations of each attack and the refusal direction on HarmBench, using cosine similarity. If the similarity is higher, the model is expressing the refusal feature, which should increase the probability of a refusal.
@@
The leftmost plot in Figure 4 compares the agreement across different attacks on Vicuna-13B-v1.5. Prompt-based attacks like GCG show less agreement, by reducing a model’s focus on harmful inputs and bypassing refusal by limiting its expression [34]. Ortho removes the direction entirely (indicated by the constant cosine similarity of zero), suppressing a model’s ability to refuse. Interestingly, our attack reduces the refusal expression less than the others, with cosine similarities closely aligning with the clean model. This indicates a strong presence of the refusal feature, which is surprising given that we show models jailbroken by PrisonBreak refuse at a lower rate. It does not control the refusal direction, thus functioning differently from Ortho.
TABLE VIII: Countermeasures against PrisonBreak. We report the ASR for each model- and jailbreaking-based defense. “WC” stands for weight clipping, “SD” SafeDecoding, and “IA” Intention Analysis. “Adaptive” indicates an adaptive attacker, described in the main body.
Models Model-Based Defenses Models Jailbreaking-Based Defenses None Ranger WC WC (Adaptive) None SD SD (Adaptive) IA IA (Adaptive) Vicuna-7b-v1.3 94.34 94.34 12.56 67.92 Vicuna-7b-v1.5 93.08 40.25 88.05 15.72 90.57 Llama2-7b 86.79 86.16 1.26 64.15 Llama2-7b 86.79 72.96 88.05 0 84.28 Qwen2-1.5b 80.50 81.13 15.09 78.62 Llama3-8b 86.79 11.32 81.76 7.55 91.19 Average 87.21 87.21 9.64 70.23 Average 88.89 41.51 85.95 7.76 88.68
@@
Decomposing activations. To analyze activation differences in jailbroken models, we decompose them using a dimensionality reduction technique, UMAP [68]. We draw 100 random samples from Alpaca and HarmBench, and for each attack, obtain the activation at the last token position. We run UMAP on activations collected from the 200 samples, and reduce them into two components [^2] We compare activations from the clean model and models compromised by Ortho and our attack. Empirically, these models are close to the clean one in loss space, indicating their activations do not differ drastically.
@@
The two remaining figures show visualizations for Qwen2-7B; we consider the last layer, where activations are most distinct and closest to the model’s output (see Appendix C for more layers). The clean model separates benign and harmful prompts into two distinct clusters. The harmful cluster may represent a region in the activation space where the model has learned to refuse. Prompt-based attacks like GCG produce activations that are similar for both harmful and benign inputs. This confirms that the adversarial suffix induces a distribution shift in the inputs, bypassing refusal by producing activations that differ from what the model can identify. Similarly, for Ortho, most activations from harmful inputs are near the benign ones, and align with the benign activations of the clean model. This suggests Ortho suppresses the refusal features, causing the jailbroken model to respond to all inputs. Our attack shows different behavior. The cluster for benign inputs overlaps significantly with that from the clean model, while harmful inputs form a distinct cluster. Our attack preserves a model’s utility but does not fully suppress the refusal features like Ortho, as shown in our mechanistic analysis. PrisonBreak may instead amplify activations associated with responding.
@@ 9 Potential Countermeasures
Potential countermeasures can defend against Rowhammer, jailbreaking, and bit-flips in the model. Many defenses have been proposed to mitigate Rowhammer [16, 69, 70], but increased latency, memory, and runtime overhead have prevented real-world deployment. Several works have also produced effective jailbreaking defenses [71, 72, 73, 74]; although designed primarily for prompt-based attacks, certain algorithms apply to other attack vectors, such as bit-flips. Model-based bit-flip defenses [75, 76, 77, 78] are related to our work, but many consider incompatible objectives or incur substantial computational overhead for commercial-scale language models. For these reasons, we focus on low-overhead and practically applicable jailbreaking- and model-based defenses, providing an in-depth overview of all defense paradigms in Appendix E.
@@
For model-based defenses, we examine Ranger [78], which clamps activations during inference, and quantized-weight reconstruction [79] which contains weights within bounds set by locally computed averages. Both defenses aim to mitigate the large parameter changes caused by bit-flip attacks. For Ranger, we compute bounds using 20% of Alpaca data (following the original study) and clip the outputs of each Transformer block. We adapt the defense by Li et al. [79] by clipping each layer’s weights to their original min. and max. values before each generation.
For the jailbreaking-based defenses, we evaluate SafeDecoding [74], which fine-tunes a separate “safer” version of the victim model to steer logits away from harmful completions, and Intention Analysis [80], which prompts the model to reason about the user’s intent and potential harms before generating a response. For SafeDecoding, we utilize the publicly available implementation and fine-tuned models. For Intention Analysis, we apply the exact ‘one-pass’ prompt introduced in the original work to all queries.
In Table VIII, we show that our attack remains effective against these defenses. Ranger is ineffective, yielding no reduction in ASR. This is likely because the large activations typical in language models [81] produce clamping bounds that are not exceeded, even after substantial weight value changes. Weight clipping (WC) effectively reduces ASR from 87.2% to 9.6%, comparable to clean models (9.4%). But, against an adaptive adversary (Adaptive) who selectively inflicts bit-flips while keeping parameter changes within specified layer-wise bounds, the attack regains effectiveness (70.2%). SafeDecoding (SD) is less effective, reducing ASR to 41.5%, though its impact varies with the robustness of the underlying safety-tuned models. An adaptive attacker targeting these fine-tuned models can restore ASR to 86.0%, nearly matching the undefended baseline. Intention Analysis (IA) is the most effective, reducing ASR to 7.8%, lower than that observed in clean models. The reformatted prompt significantly shifts the output distribution, rendering original bit-flips ineffective. However, an adaptive adversary can identify bit-flips that remain effective under IA’s reformatted prompts, recovering ASR to 88.7%.
@@ 10 Conclusion
Our work exposes a new vulnerability in commercial-scale language models: the ability of these models to refuse answering harmful queries, a feature learned through safety alignment mechanisms, can be broken by manipulating only a few parameters out of billions. This finding is surprising and in stark contrast to prior work [34, 33], which requires altering millions of parameters to remove the refusal ability. We introduce PrisonBreak, a new attack that exploits this vulnerability. The adversary induces targeted bit-flips at memory locations storing critical parameters to jailbreak language models at runtime. We also present an efficient offline procedure for identifying target bits, which reduces the computations required to locate critical bits by over an order of magnitude compared to prior attacks. The attacker then leverages Rowhammer, a software-induced fault attack, to flip those target bits in memory and jailbreak the model. In our evaluation of 10 open-source language models, we show that the attacker can PrisonBreak these models with as few as 5 and no more than 25 bit-flips. Our jailbroken models generate high-quality responses to harmful queries with an attack success of 80–98% while preserving their utility. We also demonstrate that an attacker can jailbreak target models without access to their parameters—identified bit-flips transfer between models that share architecture and pre-training. Our attack reliably jailbreaks all evaluated models even on systems targeted by prior bit-flip attacks. Moreover, our hardware-aware search algorithm can achieve successful jailbreaking on models deployed on highly Rowhammer-secure systems. We characterize the mechanisms behind our attack, and show PrisonBreak operates differently than existing jailbreaking attacks. We also discuss potential countermeasures to mitigate the risk.
@@ Appendix A Experimental Setup in Detail
We implement our attack using Python v3.10.13 and PyTorch v2.4.1, with CUDA 12.1 for GPU usage. All language models and datasets used in our work are open-source and available on HuggingFace [55] or their respective repositories. Information regarding the DRAM modules used in our study is available from the original work [39]. We run the PrisonBreak evaluation on a machine with an Intel Xeon Processor with 48 cores, 64 GB of memory, and 8 Nvidia A40 GPUs. For our fault injection simulations, we use an Intel Xeon Processor with 40 cores and 128GB of memory.
Figure 5: The ASR and Acc. across proxy dataset size, number of critical weights, utility score threshold, and number of bit-flips for Vicuna-7b-v1.3.
Appendix B Impact of Attack Configurations
Here, we vary the attack configuration choices for an additional model: Vicuna-7b-v1.3. Results are presented in Figure 5. For the proxy dataset size, using just 5 samples is sufficient to achieve a high ASR (87.4%). We also find that increasing the number of candidates from 100$\rightarrow$150 lowers ASR from 94.34$\rightarrow$91.8%. These results differ from §6.5, where we found that a dataset size of $5$ is ineffective and setting $k=150$ substantially improved ASR on Llama2-7b. It is likely that because lesser-aligned models like Vicuna-7b-v1.3 are easier to jailbreak, PrisonBreak requires less data and fewer candidates. We observe largely similar results for the utility score threshold, but find the most aggressive value ($0.001$) less effective, as increasing it to $0.01$ raises ASR from 85.5$\rightarrow$94.3with negligible impact on Acc. We find that the tighter threshold required the analysis of substantially more candidates, meaning the selected bit-flips minimized the jailbreaking objective less effectively. Like §6.5, a manual analysis of the harmful generations finds that higher thresholds ($>0.01$) result in lower fidelity outputs regardless of the reported Acc. on tinyBenchmarks. Finally, the rightmost figure shows that only 5 bit-flips are needed to achieve the best ASR, after which it saturates, supporting our finding in §6.2 that less-aligned models require fewer bit-flips to jailbreak.
Appendix C Activation Decomposition
@@
Figure 6 shows the decomposed activations from Alpaca and HarmBench for three additional layers (4, 12, and 20) using the procedure detailed in §8.2. We run UMAP independently for each layer; hence, the orientation of clusters across layers is arbitrary and not comparable. In earlier layers, we find that benign and harmful activations are not clustered apart in the clean model, as it is likely focusing on lower-level information [67] (e.g., syntax). Ortho and our attack produce activations that are virtually identical to the clean model, because the parameter perturbations are minimal (in our case, we do not flip bits before or in layer 4). GCG, however, has already formed a distinct cluster, showing that the adversarial suffix has an immediate impact. In the middle layers, we see clusters forming between benign and harmful activations for all methods but GCG. Interestingly, the harmful activations from PrisonBreak and Ortho are clustered tightly together but separate from the clean model; the attack’s act on activations similarly here, despite forming distinct clusters in later layers. By layer 20, we observe a similar clustering to layer 12 except that the harmful activations from our attack and Ortho have separated. Many bit flips made by PrisonBreak are concentrated in these middle–final layers, evidently inducing a unique shift in the activation space.
Appendix D Results on tinyBenchmarks
@@
TABLE IX: Full tinyBenchmarks results. The accuracy change from PrisonBreak on each tinyBenchmark task for all models.
Models tinyBenchmarks Tasks Arc GSM8k Hella- Swag MMLU Truth- FulQA Wino- Grande Vicuna-7B-v1.3 -0.58 +0.38 -1.16 -0.54 -1.98 -4.08 Vicuna-7B-v1.5 +0.03 -3.04 -0.97 -0.61 -7.92 -2.57 Vicuna-13B-v1.5 -1.34 -4.60 -4.98 -0.84 -5.90 -6.87 Llama2-7B +0.88 -4.54 -4.51 -1.77 -4.14 +2.58 Llama2-13B -5.68 -12.03 -4.53 -1.14 -0.53 -1.09 Llama3-8B -0.55 -2.40 -0.50 +0.50 +0.07 +3.22 Llama3-70B -6.25 -7.39 -0.12 -0.94 -1.84 +1.16 Tulu3-8B -0.14 -7.39 +0.87 +0.82 -2.46 +2.59 Qwen2-1.5B -2.26 -11.44 -1.42 -1.25 -2.88 -5.32 Qwen2-7B +0.49 -3.67 -0.42 -2.47 +0.08 -0.04 Average -1.54 -5.61 -1.77 -0.82 -2.75 -1.04
@@
Table IX shows our full evaluation results for PrisonBreak on tinyBenchmarks. Notably, the change in performance differs across tasks. Models on average perform the worst on GSM8k—a benchmark consisting of grade school math problems—with an accuracy reduction of 5.61%. This may arise from the lack of these problems in the dataset we use to compute the utility score. The second most affected task is TruthfulQA with an average accuracy reduction of 2.75%. This fact is less surprising as jailbroken models are more likely to be deceptive [33, 34]. Other tasks exhibit smaller drops in accuracy; in general, we find that PrisonBreak preserves utility on benign tasks.